Solve the system by elimination method
x=9y+8
3x+y=-60

Answer by mananth(16946)

1 x -9 y = 8 .............1

3 x + 1 y = -60 .............2
Eliminate y
multiply (1)by 1
Multiply (2) by 9
1 x -9 y = 8
27 x + 9 y = -540
Add the two equations
28 x = -532
/ 28
x = -19
plug value of x in (1)
1 x -9 y = 8
-19 -9 y = 8
-9 y = 8 + 19
-9 y = 27
y = -3 1
x= -19
y= -3
m.ananth@hotmail.ca